# Efficient and Comprehensive Feature Extraction in Large Vision-Language Model for Clinical Pathology Analysis

### v1

@@ Abstract
Pathological diagnosis is vital for determining disease characteristics, guiding treatment, and assessing prognosis, relying heavily on detailed, multi-scale analysis of high-resolution whole slide images (WSI). However, traditional pure vision models face challenges of redundant feature extraction, whereas existing large vision-language models (LVLMs) are limited by input resolution constraints, hindering their efficiency and accuracy. To overcome these issues, we propose two innovative strategies: the mixed task-guided feature enhancement, which directs feature extraction toward lesion-related details across scales, and the prompt-guided detail feature completion, which integrates coarse- and fine-grained features from WSI based on specific prompts without compromising inference speed. Leveraging a comprehensive dataset of 490,000 samples from diverse pathology tasks—including cancer detection, grading, vascular and neural invasion identification, and so on—we trained the pathology-specialized LVLM, OmniPath. Extensive experiments demonstrate that this model significantly outperforms existing methods in diagnostic accuracy and efficiency, offering an interactive, clinically aligned approach for auxiliary diagnosis in a wide range of pathology applications.
Figure 1: From top to bottom are examples of our OmniPath, a vision-language model optimized for pathology, applied to referring expression detection, segmentation, and visual question answering. Notably, in the first example, OmniPath is tasked with detecting cancer cell nuclei within blood vessels. Results show that OmniPath accurately identifies most nuclei within vessels without mistakenly detecting any outside, demonstrating its capability to understand pathological concepts and reason effectively.
@@ 1 Introduction
To make LVLM truly applicable to clinical pathology for auxiliary diagnosis, we gathered visual instruction-following data for multiple organs and tasks from several institutions, based on diagnostic items in actual pathology reports. This data trains the model to meet the diverse clinical needs of pathologists through human-computer interaction. These tasks include cancer region detection and segmentation, cancer grading and classification, identification of vascular and neural invasion, and lymph node metastasis detection, among others, with each task containing meticulously compiled datasets ranging from hundreds to thousands of instruction-following data samples. Furthermore, to strengthen the model’s understanding of foundational pathology concepts, we collected training data for fundamental tasks such as nucleus detection and classification, vascular and neural detection, lymph node detection, and tumor-infiltrating lymphocyte identification. Additionally, we integrated pathology image-text data from publicly available online resources, including the PubMed database [17, 40], pathology textbooks and atlases [20, 40], The Cancer Genome Atlas (TCGA) [39, 9], Twitter posts [21, 41], and educational histopathology videos on YouTube [22, 36]. This integration yielded a comprehensive dataset covering 20 organs with approximately 490,000 training samples. Leveraging our efficient and comprehensive feature extraction scheme and this extensive dataset, we trained OmniPath, a specialized LVLM for pathology, capable of providing comprehensive pathology auxiliary diagnostic services through human-computer interaction. Extensive experimental results demonstrate that OmniPath outperforms existing pathology LVLMs across a range of diagnostic tasks, better aligning with the actual needs of clinical practice.
@@
  • We integrated multi-source data to construct a pathology visual instruction tuning dataset covering 20 organs with approximately 490,000 training samples. Based on this dataset, we trained the OmniPath pathology large vision-language model. Extensive experimental validation shows that OmniPath surpasses existing pathology large vision-language models across various diagnostic tasks, aligning more closely with clinical pathology diagnostic needs through an interactive paradigm.
@@ 3.1 Preliminaries
Figure 3: Overview of the proposed OmniPath. Left: the architecture of OmniPath with feature enhancement and completion strategy. Right: the detailed structure of the multi-scale feature fusion vision encoder.
@@ 3.2 Decision-Dependent Image Tokens Analysis
To further investigate the image feature patterns that LVLM relies on during the answer generation and decision-making processes, and to optimize the model to focus more effectively on task-relevant features, thereby improving accuracy in responding to human queries, we conducted an attention pattern analysis on existing medical LVLMs. Specifically, we selected two representative models, LLaVA-Med [27] and Quilt-LLaVA [36], using a unified prompt, “What cancer subtype is shown in this image?” to guide the models in performing cancer subtype identification on pathology slides. Our proposed OmniPath model was included as a comparison. Visualization of the relevant analysis results is shown in Figure 2.
@@
In contrast, in the optimized OmniPath model, the heatmap of $\Psi_{(N+M),e_{v}}$ shows that key image tokens are concentrated in the cancerous regions, indicating that these areas contribute more information to $\mathcal{M}$, aiding in accurate cancer subtype identification. From the heatmap of $\Psi_{e_{v}^{k},e_{v}}$, it can be seen that although the key token maintains high attention values across all preceding image tokens, its focus on tokens within the cancerous regions is significantly higher. This demonstrates that OmniPath can more precisely capture critical diagnostic features in pathology WSIs, thus achieving a better response to user instructions. The next section will elaborate on the optimization strategies employed in OmniPath.
@@ 4 Method
To address the identified limitations of existing LVLMs, specifically their tendency to over-rely on key tokens and inability to comprehensively capture multi-scale pathological features, we propose a novel framework that enhances both feature extraction precision and efficiency. Our approach consists of two complementary strategies: the mixed task-guided feature enhancement (MTGFE) and the prompt-guided detail feature completion (PGDFC). These strategies work in concert to improve the model’s capability in pathological image analysis by targeting the core challenges revealed in our previous analysis while maintaining computational efficiency and diagnostic accuracy. OmniPath trained with these two strategies is shown in Figure 3. Below, we elaborate on these strategies and their implementation.
@@ 5 Experiments
In this section, we first provide the implementation details of OmniPath, followed by comparative results across multiple tasks. Finally, we conduct ablation studies on the key components of OmniPath.
@@ 5.1 Implementation Details
Based on the pretrained LLaVA-1.5-13B [30], we constructed OmniPath by replacing its CLIP ViT-L 336px [35] visual encoder with SigLIP ViT-SO 384px [53] and integrating UNI [11] as an auxiliary vision encoder. The projector utilizes a two-layer MLP with GELU activation. The mask encoder is implemented with ResNet-50 [19], while the mask decoder follows SAM’s [24] decoder architecture but directly uses the LVLM’s vision encoder to replace SAM’s original image encoder for feature extraction. During training, all modules of OmniPath participate in end-to-end training with no parameter freezing. We created a multitask dataset encompassing 20 organs and approximately 490,000 samples for model training (see supplementary material for detailed data sources and construction methods). Unlike the two-stage training strategy commonly adopted by existing LVLMs, OmniPath requires only a single-stage fine-tuning: trained for 2 epochs on 8 NVIDIA A800 GPUs using the AdamW optimizer, with a learning rate of 2e-5 and a global batch size of 128. We set $S=8$, and compare OmniPath with LLaVA-1.5 [30], LLaVA-Med [27], Quilt-LLaVA [36], and PA-LLaVA [13].
@@ 5.2 Comparison on Pathology Diagnostic Tasks
It can be observed that OmniPath achieves the best performance across all patch-level and slide-level pathological diagnosis tasks. For most cancer subtype classification and grading tasks, OmniPath achieves accuracy rates exceeding 70%, and in many cases, surpassing 90%. In contrast, the accuracy rates of other models are generally below 70%. This demonstrates that OmniPath is more suitable for clinical applications to assist pathologists in diagnosis. Furthermore, OmniPath also exhibits strong recognition capabilities for features such as microvascular invasion, neural invasion, and tumor-infiltrating lymphocytes. This significantly reduces the extensive effort required by pathologists to meticulously examine detailed pathological lesions during slide review.
@@ 5.3 Comparison on Zero-Shot Classification Tasks
To evaluate the clinical generalization capability of OmniPath, we employed a zero-shot classification paradigm, testing on several widely-used academic pathology datasets that were not included in the training set. The evaluation covered two levels: patch-level tasks using the CCRCC [6], MHIST [47], and NCT-CRC [32] datasets, and slide-level tasks using the PANDA [7], DHMC [48], and CAMELYON17 [29] datasets. Using a closed-ended question-answering approach, the model was required to classify images into predefined categories specific to each dataset. The performance comparison of all models on these zero-shot test sets is presented in Table 4.
It is shown that OmniPath consistently outperforms other models across both patch-level and slide-level datasets in zero-shot classification tasks, highlighting its strong generalization ability in pathological image analysis. Notably, on the PANDA and CAMELYON17 slide-level datasets, OmniPath achieved the highest accuracy rates of 79.15% and 59.33%, respectively, which significantly surpasses the performance of other models. This superior performance in zero-shot classification indicates OmniPath’s robustness in handling diverse pathological image data and reinforces its potential for clinical applications where labeled training data may be limited.
@@ 5.4 Detection and Segmentation Performance
In pathological diagnosis, detection and segmentation tasks are as critical as classification tasks. For instance, assessing microvascular invasion requires counting cancer cell nuclei within blood vessels, which relies on accurate referring detection. Given that existing pathological LVLMs lack detection and segmentation capabilities, this part presents only OmniPath’s performance on these tasks. The detection tasks cover various cancer region identifications, including HCC, ICC, RCC, glioblastoma (GBM), lung adenocarcinoma (LUAD), and bladder cancer (BC). Additionally, it involves detecting tissue structures such as lymph nodes (LD), vessels (VD), and nerves (ND), as well as cell nuclei detection in datasets like MoNuSeg [25] (without categories), NuCLS [3], and PanNuke [18] (with categories). The segmentation tasks not only include the segmentation of cancer regions covered by the detection tasks but also nerve segmentation (NS), nerve invasion region segmentation (NIS), and lymph node metastasis segmentation (LNMS). Detection tasks are evaluated using F1-score and IoU, while segmentation tasks are assessed using the IoU and Dice coefficient. Table 3 and Table 5 show the detection and segmentation performance, respectively.
While OmniPath’s performance may not surpass specialized smaller models on certain tasks, it offers a unique advantage in inferential capabilities. As shown in Figure 1, OmniPath can accurately detect cancer cell nuclei within blood vessels according to instructions—an ability that current specialized small models find challenging to achieve.
@@ 5.5 Ablation Study
The ablation study in Table 6 underscores the effectiveness of the PGDFC strategy in enhancing OmniPath’s performance across WSI diagnostic tasks. Three configurations were tested: (1) without PGDFC, (2) replacing the top-S elements in $\Psi_{(N+M),e_{v}}$ with random selection of $\hat{x}_{v}$, and (3) with the designed PGDFC strategy. Removing PGDFC resulted in an average accuracy drop of 21.1%, while random selection led to a decrease of 12.7% compared to using PGDFC. Notably, for tasks like HCC-S and ICC-G, PGDFC boosted accuracy significantly, demonstrating its ability to enhance model focus on essential features, which is critical for accurate diagnosis across complex pathology tasks. In certain tasks, such as HCC-G and STAD-L, random selection results in performance that is even lower than without PGDFC, indicating that incorrectly supplementing detailed features can also impair model performance.
@@ 6 Conclusion
This paper presents OmniPath, a pathology-focused LVLM shaped by two key strategies addressing limitations in existing models. The mixed task-guided feature enhancement strategy ensures precise lesion-specific feature extraction, while the prompt-guided detail feature completion strategy combines global context with fine-grained detail to meet clinical needs. Together, these approaches enable comprehensive and balanced feature extraction, validated across diverse pathology tasks, positioning OmniPath as a transformative tool in digital pathology.
Limitations and Future Work. OmniPath faces three main limitations: First, the model lacks sufficient depth in medical and pathological expertise due to training data primarily consisting of image-caption pairs, with limited integration of advanced pathology knowledge and literature. To address this, we plan to enhance the model’s knowledge base using RAG techniques. Second, its performance in zero-shot classification tasks remains suboptimal. We will introduce a multi-agent framework to enable specialized agents to assist with more accurate diagnoses in complex cases. Finally, the model’s reasoning capability is not yet sufficient for independent diagnosis. To resolve this, we will collect pathologists’ diagnostic process data and integrate it with larger-scale LLMs, aiming to enhance reasoning and achieve autonomous diagnosis, ultimately reducing physicians’ workload.

### v2

@@ Abstract
Pathological diagnosis is vital for determining disease characteristics, guiding treatment, and assessing prognosis, relying heavily on detailed, multi-scale analysis of high-resolution whole slide images (WSI). However, traditional pure vision models face challenges of redundant feature extraction, whereas existing large vision-language models (LVLMs) are limited by input resolution constraints, hindering their efficiency and accuracy. To overcome these issues, we propose two innovative strategies: the mixed task-guided feature enhancement, which directs feature extraction toward lesion-related details across scales, and the prompt-guided detail feature completion, which integrates coarse- and fine-grained features from WSI based on specific prompts without compromising inference speed. Leveraging a comprehensive dataset of 490,000 samples from diverse pathology tasks—including cancer detection, grading, vascular and neural invasion identification, and so on—we trained the pathology-specialized LVLM, OmniPT. Extensive experiments demonstrate that this model significantly outperforms existing methods in diagnostic accuracy and efficiency, offering an interactive, clinically aligned approach for auxiliary diagnosis in a wide range of pathology applications.
Figure 1: From top to bottom are examples of our OmniPT, a vision-language model optimized for pathology, applied to referring expression detection, segmentation, and visual question answering. Notably, in the first example, OmniPT is tasked with detecting cancer cell nuclei within blood vessels. Results show that OmniPT accurately identifies most nuclei within vessels without mistakenly detecting any outside, demonstrating its capability to understand pathological concepts and reason effectively.
@@ 1 Introduction
To make LVLM truly applicable to clinical pathology for auxiliary diagnosis, we gathered visual instruction-following data for multiple organs and tasks from several institutions, based on diagnostic items in actual pathology reports. This data trains the model to meet the diverse clinical needs of pathologists through human-computer interaction. These tasks include cancer region detection and segmentation, cancer grading and classification, identification of vascular and neural invasion, and lymph node metastasis detection, among others, with each task containing meticulously compiled datasets ranging from hundreds to thousands of instruction-following data samples. Furthermore, to strengthen the model’s understanding of foundational pathology concepts, we collected training data for fundamental tasks such as nucleus detection and classification, vascular and neural detection, lymph node detection, and tumor-infiltrating lymphocyte identification. Additionally, we integrated pathology image-text data from publicly available online resources, including the PubMed database [17, 40], pathology textbooks and atlases [20, 40], The Cancer Genome Atlas (TCGA) [39, 9], Twitter posts [21, 41], and educational histopathology videos on YouTube [22, 36]. This integration yielded a comprehensive dataset covering 20 organs with approximately 490,000 training samples. Leveraging our efficient and comprehensive feature extraction scheme and this extensive dataset, we trained OmniPT, a specialized LVLM for pathology, capable of providing comprehensive pathology auxiliary diagnostic services through human-computer interaction. Extensive experimental results demonstrate that OmniPT outperforms existing pathology LVLMs across a range of diagnostic tasks, better aligning with the actual needs of clinical practice.
@@
  • We integrated multi-source data to construct a pathology visual instruction tuning dataset covering 20 organs with approximately 490,000 training samples. Based on this dataset, we trained the OmniPT pathology large vision-language model. Extensive experimental validation shows that OmniPT surpasses existing pathology large vision-language models across various diagnostic tasks, aligning more closely with clinical pathology diagnostic needs through an interactive paradigm.
@@ 3.1 Preliminaries
Figure 3: Overview of the proposed OmniPT. Left: the architecture of OmniPT with feature enhancement and completion strategy. Right: the detailed structure of the multi-scale feature fusion vision encoder.
@@ 3.2 Decision-Dependent Image Tokens Analysis
To further investigate the image feature patterns that LVLM relies on during the answer generation and decision-making processes, and to optimize the model to focus more effectively on task-relevant features, thereby improving accuracy in responding to human queries, we conducted an attention pattern analysis on existing medical LVLMs. Specifically, we selected two representative models, LLaVA-Med [27] and Quilt-LLaVA [36], using a unified prompt, “What cancer subtype is shown in this image?” to guide the models in performing cancer subtype identification on pathology slides. Our proposed OmniPT model was included as a comparison. Visualization of the relevant analysis results is shown in Figure 2.
@@
In contrast, in the optimized OmniPT model, the heatmap of $\Psi_{(N+M),e_{v}}$ shows that key image tokens are concentrated in the cancerous regions, indicating that these areas contribute more information to $\mathcal{M}$, aiding in accurate cancer subtype identification. From the heatmap of $\Psi_{e_{v}^{k},e_{v}}$, it can be seen that although the key token maintains high attention values across all preceding image tokens, its focus on tokens within the cancerous regions is significantly higher. This demonstrates that OmniPT can more precisely capture critical diagnostic features in pathology WSIs, thus achieving a better response to user instructions. The next section will elaborate on the optimization strategies employed in OmniPT.
@@ 4 Method
To address the identified limitations of existing LVLMs, specifically their tendency to over-rely on key tokens and inability to comprehensively capture multi-scale pathological features, we propose a novel framework that enhances both feature extraction precision and efficiency. Our approach consists of two complementary strategies: the mixed task-guided feature enhancement (MTGFE) and the prompt-guided detail feature completion (PGDFC). These strategies work in concert to improve the model’s capability in pathological image analysis by targeting the core challenges revealed in our previous analysis while maintaining computational efficiency and diagnostic accuracy. OmniPT trained with these two strategies is shown in Figure 3. Below, we elaborate on these strategies and their implementation.
@@ 5 Experiments
In this section, we first provide the implementation details of OmniPT, followed by comparative results across multiple tasks. Finally, we conduct ablation studies on the key components of OmniPT.
@@ 5.1 Implementation Details
Based on the pretrained LLaVA-1.5-13B [30], we constructed OmniPT by replacing its CLIP ViT-L 336px [35] visual encoder with SigLIP ViT-SO 384px [53] and integrating UNI [11] as an auxiliary vision encoder. The projector utilizes a two-layer MLP with GELU activation. The mask encoder is implemented with ResNet-50 [19], while the mask decoder follows SAM’s [24] decoder architecture but directly uses the LVLM’s vision encoder to replace SAM’s original image encoder for feature extraction. During training, all modules of OmniPT participate in end-to-end training with no parameter freezing. We created a multitask dataset encompassing 20 organs and approximately 490,000 samples for model training (see supplementary material for detailed data sources and construction methods). Unlike the two-stage training strategy commonly adopted by existing LVLMs, OmniPT requires only a single-stage fine-tuning: trained for 2 epochs on 8 NVIDIA A800 GPUs using the AdamW optimizer, with a learning rate of 2e-5 and a global batch size of 128. We set $S=8$, and compare OmniPT with LLaVA-1.5 [30], LLaVA-Med [27], Quilt-LLaVA [36], and PA-LLaVA [13].
@@ 5.2 Comparison on Pathology Diagnostic Tasks
It can be observed that OmniPT achieves the best performance across all patch-level and slide-level pathological diagnosis tasks. For most cancer subtype classification and grading tasks, OmniPT achieves accuracy rates exceeding 70%, and in many cases, surpassing 90%. In contrast, the accuracy rates of other models are generally below 70%. This demonstrates that OmniPT is more suitable for clinical applications to assist pathologists in diagnosis. Furthermore, OmniPT also exhibits strong recognition capabilities for features such as microvascular invasion, neural invasion, and tumor-infiltrating lymphocytes. This significantly reduces the extensive effort required by pathologists to meticulously examine detailed pathological lesions during slide review.
@@ 5.3 Comparison on Zero-Shot Classification Tasks
To evaluate the clinical generalization capability of OmniPT, we employed a zero-shot classification paradigm, testing on several widely-used academic pathology datasets that were not included in the training set. The evaluation covered two levels: patch-level tasks using the CCRCC [6], MHIST [47], and NCT-CRC [32] datasets, and slide-level tasks using the PANDA [7], DHMC [48], and CAMELYON17 [29] datasets. Using a closed-ended question-answering approach, the model was required to classify images into predefined categories specific to each dataset. The performance comparison of all models on these zero-shot test sets is presented in Table 4.
It is shown that OmniPT consistently outperforms other models across both patch-level and slide-level datasets in zero-shot classification tasks, highlighting its strong generalization ability in pathological image analysis. Notably, on the PANDA and CAMELYON17 slide-level datasets, OmniPT achieved the highest accuracy rates of 79.15% and 59.33%, respectively, which significantly surpasses the performance of other models. This superior performance in zero-shot classification indicates OmniPT’s robustness in handling diverse pathological image data and reinforces its potential for clinical applications where labeled training data may be limited.
@@ 5.4 Detection and Segmentation Performance
In pathological diagnosis, detection and segmentation tasks are as critical as classification tasks. For instance, assessing microvascular invasion requires counting cancer cell nuclei within blood vessels, which relies on accurate referring detection. Given that existing pathological LVLMs lack detection and segmentation capabilities, this part presents only OmniPT’s performance on these tasks. The detection tasks cover various cancer region identifications, including HCC, ICC, RCC, glioblastoma (GBM), lung adenocarcinoma (LUAD), and bladder cancer (BC). Additionally, it involves detecting tissue structures such as lymph nodes (LD), vessels (VD), and nerves (ND), as well as cell nuclei detection in datasets like MoNuSeg [25] (without categories), NuCLS [3], and PanNuke [18] (with categories). The segmentation tasks not only include the segmentation of cancer regions covered by the detection tasks but also nerve segmentation (NS), nerve invasion region segmentation (NIS), and lymph node metastasis segmentation (LNMS). Detection tasks are evaluated using F1-score and IoU, while segmentation tasks are assessed using the IoU and Dice coefficient. Table 3 and Table 5 show the detection and segmentation performance, respectively.
While OmniPT’s performance may not surpass specialized smaller models on certain tasks, it offers a unique advantage in inferential capabilities. As shown in Figure 1, OmniPT can accurately detect cancer cell nuclei within blood vessels according to instructions—an ability that current specialized small models find challenging to achieve.
@@ 5.5 Ablation Study
The ablation study in Table 6 underscores the effectiveness of the PGDFC strategy in enhancing OmniPT’s performance across WSI diagnostic tasks. Three configurations were tested: (1) without PGDFC, (2) replacing the top-S elements in $\Psi_{(N+M),e_{v}}$ with random selection of $\hat{x}_{v}$, and (3) with the designed PGDFC strategy. Removing PGDFC resulted in an average accuracy drop of 21.1%, while random selection led to a decrease of 12.7% compared to using PGDFC. Notably, for tasks like HCC-S and ICC-G, PGDFC boosted accuracy significantly, demonstrating its ability to enhance model focus on essential features, which is critical for accurate diagnosis across complex pathology tasks. In certain tasks, such as HCC-G and STAD-L, random selection results in performance that is even lower than without PGDFC, indicating that incorrectly supplementing detailed features can also impair model performance.
@@ 6 Conclusion
This paper presents OmniPT, a pathology-focused LVLM shaped by two key strategies addressing limitations in existing models. The mixed task-guided feature enhancement strategy ensures precise lesion-specific feature extraction, while the prompt-guided detail feature completion strategy combines global context with fine-grained detail to meet clinical needs. Together, these approaches enable comprehensive and balanced feature extraction, validated across diverse pathology tasks, positioning OmniPT as a transformative tool in digital pathology.
Limitations and Future Work. OmniPT faces three main limitations: First, the model lacks sufficient depth in medical and pathological expertise due to training data primarily consisting of image-caption pairs, with limited integration of advanced pathology knowledge and literature. To address this, we plan to enhance the model’s knowledge base using RAG techniques. Second, its performance in zero-shot classification tasks remains suboptimal. We will introduce a multi-agent framework to enable specialized agents to assist with more accurate diagnoses in complex cases. Finally, the model’s reasoning capability is not yet sufficient for independent diagnosis. To resolve this, we will collect pathologists’ diagnostic process data and integrate it with larger-scale LLMs, aiming to enhance reasoning and achieve autonomous diagnosis, ultimately reducing physicians’ workload.